

********************************************************************
UARS WEEKLY STATUS REPORT - 24 DEC 1999
********************************************************************
***********
SOLSTICE
***********
All SOLSTICE science objectives are being met during this period of reduced
tape recorder operations. We've been able to retrieve about 60% of the
planned experiements this week with SOLSTICE operating normally, collecting
solar data on all available orbital tracking periods and stellar calibration
observations, whenever possible.
Happy Winter Solstice!
***********
SUSIM
***********
SUSIM gathered a minimum set of daily solar scans this past week. The
fraction of data that the instrument gathered that we actually received
in telemetry is improving because the downlink pass times are better
known farther in advance.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM team wishes everyone Happy Holidays!
Daily activity plans for the PEM instrument were successfully generated
for the first two weeks of 2000. PEM health and safety data monitoring, which
was reconfigured for day light operation, are being reviewed to determine
if they can be modified to account for telemetry dropouts caused by the tape
recorder failure.
The PEM spectrometers continue to be active under PMON control*. While
under PMON control, the PEM spectrometers computed energy deposition from
electrons and protons measured in-situ by the particle spectrometers.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The schedule for instrument operations for
the yaw period 6 January to 15 February was distributed via
e-mail.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 10 December, 1999 (Orbit 45080, GMT Day 345) through 17
December 1999 (Orbit 45184, GMT Day 351).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 346/1245 45102 On
WINDII 349/0054 45140 On
The beta angle ranged between the angles of 6.1 degrees to 26.8 degrees for
this report period. The beta angle is now increasing toward a maximum angle of
+34.0 degrees on 23 December 1999 (DOY 357).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.3 degrees C. Additional battery performance data for each
battery on 10 December 1999 and 17 December 1999 are provided in Appendix D.
The clock error ranged from 12.5 to -11.7 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 00
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 03
% HALOE = 02
% HRDI = 01
% WINDII = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 03
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
% Transponder A: 99/320 - 2287.496800
% Transponder B: 99/334 - 2287.499196
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 99/350 show the observatory data loss to be 342 hours, 08
minutes, 08.729 seconds (an increase of 69 hours 17 minutes 00.192 seconds
since the last report period). This is a 0.4741 percent data loss which equals
a 99.5259 percent data capture for the mission.
The increase of 69 hours 17 minutes 00.192 seconds occurred over seven (7)
days:
- 11 hours 40 minutes 07.840 seconds on Day 344 due to a S/C TR problem.
- 06 hours 41 minutes 49.952 seconds on Day 345 due to a S/C TR problem.
- 10 hours 34 minutes 28.896 seconds on Day 346 due to a S/C TR problem.
- 10 hours 34 minutes 50.144 seconds on Day 347 due to a S/C TR problem.
- 11 hours 27 minutes 41.440 seconds on Day 348 due to a S/C TR problem.
- 10 hours 46 minutes 29.184 seconds on Day 349 due to a S/C TR problem.
- 07 hours 31 minutes 32.736 seconds on Day 350 due to a S/C TR problem.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
A positive 0.4 degree roll maneuver was successfully completed for WINDII.
They requested this temporary attitude maneuver from December 15 thru December
23, 1999 for science observations. No other science systems were impacted
during this period.
A reoccuring timing error on the PEM instrument has been isolated by the PEM
team to be related to a heater being cycled just prior to their activation
sequence. This triggers a false out-of-limits error but has no impact to the
operation of the PEM instrument. PEM has given the FOT the directive to ignore
the error for now.
YEAR 2000 OPERABILITY SUPPORT
All areas identified to or by the Flight Operations Team for Y2K testing have
been completed. There are no other areas identified at this time that have not
been Y2K certified.
MEETINGS
The FOT did not attend any meetings during this report period.
FUTURE SIGNIFICANT EVENTS
No specific date was agreed upon for UARS Safehold configuration discussions.
Instead, the Supporting Engineers involved will be contacted every two weeks to
determine if either the Flight Operations Team or Supporting Engineers have any
new developments that should be discussed in a group.
UARS has tested and prepared for the Year 2000 event. Special considerations
have been made in preparation for a non nominal 1999 to 2000 transition.
The next UARS Yaw slew is scheduled for January 6, 2000. This maneuver will be
a Forward to Reverse Flight maneuver (Northern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 17 December 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
45102 346/1245 99-414 HALOE Power ON
45104 346/1432 99-415 HALOE Power ON
45140 349/0054 99-416 WINDII Power ON
45149 349/1523 99-417 ACS ROLL MANUEVER
45162 350/1301 99-418 OBC Clk Adj fr 0.00 to 20.6 msec/day
45183 351/2218 99-419 HRDI Disable Etalon Tuning Temp
45184 351/2320 99-420 OBC Clk Adj fr 20.6 to 0.00 msec/day
______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 17 December 1999
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
No Out-Of-Limits to report this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
____________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 17 December 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
45162 350/13:15 (16 Dec 99) from 0.00 to 20.6 msec/day
45184 351/23:25 (17 Dec 99) from 20.6 to 0.00 msec/day
______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 17 December 1999
10 Dec 1999 - GMT Day 344 Beta = 6.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- ------- -------- ---- ------ ------
1 -44.8/-72.8 -3.68/-4.62 +0.0/+0.0 16.2 0.000 0.0
2 +16.8/-44.8 +5.15/+3.87 +30.4/-13.2 25.8 1.099 24.8
3 +39.2/-11.2 +3.87/+2.60 +29.6/-13.2 25.8 1.087 24.5
17 Dec 1999 - GMT Day 351 Beta = 26.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- --------- ---------- --------- ---- ------ -----
1 -44.8/-67.2 -2.73/-3.68 +0.0/+0.0 16.2 0.000 0.0
2 +39.2/-44.8 +7.09/+5.47 +31.2/-15.2 24.5 1.030 21.1
3 +33.6/-16.8 +5.79/+4.51 +30.8/-16.4 24.3 1.023 21.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 17 December 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-089 FOT 45095 346/0123 14170 Generic Late Acquisition
99-090 FOT 45130 348/0829Z 21695 IP Broadcast problem/14
mins RTD lost
99-091 FOT 45179 351/1453 N/A TAC 4 Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-090 FOT 45130 348/0829Z 21695 IP Broadcast problem/14
mins RTD lost
99-091 FOT 45179 351/1453 N/A TAC 4 Crashed
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiving 512 PB
99-089 FOT 45095 346/0123 14170 Generic Late Acquisition
SEVNTFW99/51